OpenStudy (anonymous):

do i distribute the 70 in this problem sqrt 70 (sqrt14)(sqrt22)

OpenStudy (anonymous):

you can't simplify sqrt 70

OpenStudy (anonymous):

do i just multiply them all together

OpenStudy (anonymous):

try it

OpenStudy (anonymous):

the numbers to big

OpenStudy (anonymous):

its sqrt of 21560

OpenStudy (anonymous):

yes, find factors of it that are perfect cubes, such as 4

OpenStudy (anonymous):

49 is a factor

OpenStudy (anonymous):

can be simplified into 7sqrt440

OpenStudy (anonymous):

14sqrt 10

OpenStudy (anonymous):

nvm thats wrong

OpenStudy (anonymous):

14sqrt110

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes

OpenStudy (phi):

when you have \[ \sqrt{70}\cdot\sqrt{14}\cdot\sqrt{22} \] you can put the whole thing under one square root sign: \[ \sqrt{70 \cdot 14\cdot 22} \] to simplify this mess, break each number into its prime factors. (so you need to know what numbers are prime) 70= 2*5*7 (to find these, divide by 2 to get 35, then divide 35 by 5 to get 7) 14= 2*7 22= 2*11 now look for pairs (those we can pull out of the square root) 2*5*7*2*7*2*11 rearrange as 2*2*7*7 * 5*2*11 that means we can pull out square root of 2*2*7*7 to get 2*7 the answer is \[ 2\cdot 7\sqrt{2\cdot5\cdot11} = 14\sqrt{110}\]